OpenStudy (anonymous):

When rolling 2 dice, one red and one green, calculate the probabilities given that A is the event roll a total of 5, B is the event the green outcome is 2 more than red outcome, and c is the event the total is greater than 7. Find the probability of: P(A or C)

OpenStudy (anonymous):

"calculate the probabilities given that A is the event roll a total of 5, B is the event the green outcome is 2 more than red outcome, and c is the event the total is greater than 7. so the law of probability says that joint probability is found by multiplying the probabilites of each event to occur on its own right?

OpenStudy (anonymous):

This is a mutually exclusive event

OpenStudy (anonymous):

its not joint probability

OpenStudy (anonymous):

thanks I might not be the best help then... its been a long time since i've done statistics

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

I did some research and found this page: http://www.mathsisfun.com/data/probability-events-mutually-exclusive.html which informs me the operation is to add the two probabilities together. The probability of event A ocuring would be 1 + 4 2 + 3 3 + 2 4 + 1 so 4 ways we can obtain a total of 5. the total number of possible outcomes is 6^2 = 36 Then the probability that Event A will occur is 4/36 The probability of event C occurring is: the number of ways, 8, 9, 10, 11 and 12 can be made using two dice, so 6+2=8 6+3=9 6+4=10 6+5=11 6+6=12 Thus, there are 5 ways to obtain a total greater than 7 Then the probability of Event C occurring is 7/36 Thus, the probability of P(A or C) = 4/36 + 7/36

OpenStudy (anonymous):

oh ok that makes sense thanks man

OpenStudy (anonymous):

no problem, its good for me to keep up with this stuff aswell. so, win win for both of us